OpenStudy (anonymous):

Write the solution in lowest terms. 5/8 - 3/4 = 2/8 1/8 -1/8

OpenStudy (anonymous):

-1/8

OpenStudy (anonymous):

5/8 - 3/4 = 5/8 - 6/8 = -1/8